# Iterated Invariant Extended Kalman Filter (IterIEKF)

### v4

@@ IV Application of engineering interest
We propose using the IterIEKF to estimate the position, velocity, and orientation (extended pose) of the hook of a crane transporting a load, equipped with an IMU. Mounting an IMU on a crane hook for real-time sensor data transmission is technically very feasible nowadays [35], and may open the door to new automation capabilities since it allows for feedback. Leveraging the Lie group $SE_{2}(3)$ introduced by invariant filtering theory in [12], we represent the state as
@@
where $\mathbf{R}_{k}\in SO(3)$ is the rotation matrix between the IMU and inertial frames, and $\mathbf{v}_{k},\,\mathbf{p}_{k}\in\mathbb{R}^{3}$ are the IMU velocity and position in the inertial frame. The inertial frame is fixed at the crane cable attachment point with the $z$-axis oriented upward, and remains stationary. The hook and IMU frames are assumed perfectly aligned. See our preliminary conference paper [34] for more details on this application.
@@ V Conclusion
We applied the IterIEKF to the problem of estimating the extended pose of a crane hook equipped with an IMU, where mechanical information is assimilated in the form of a noise-free pseudo-measurement. The IterIEKF demonstrated the best performance, closely followed by the adapted LG-IterEKF [5], a fact explained by our present theory. The superiority of both algorithms is consistent with the theoretical guarantees we have proved, and the IEKF is outperformed in this context.
However, we recommend using the IterIEKF over the LG-IterEKF, because the IEKF possesses convergence properties (notably log-linearity of the error [12]) and consistency properties that have never been proved for the LG-EKF. In particular, the unobservability consistency properties of the IEKF [13, 19, 20] may not be inherited by the LG-IterEKF, although proving this fact is left for future research.
@@ A-A Standard case: noisy measurements
Akin to the conventional iterated EKF [2], we intend to solve (6) iteratively using the Gauss-Newton method.
Lemma 2 (from [2]).
@@
with $\mathbf{H}_{k}$ as in (10), and where the right Lie Jacobian [30] is defined by $\exp_{G}(\boldsymbol{\xi}+\boldsymbol{\delta})\approx\exp_{G}(\boldsymbol{\xi})\exp_{G}(\mathbf{\mathcal{J}}_{r}(\boldsymbol{\xi})\boldsymbol{\delta})$ neglecting terms of order $\|\boldsymbol{\delta}\|^{2}$, see (30) for an expression.
@@
with $\mathbf{H}_{k}=\mathbf{H}_{k}^{0}$ and $\mathbf{K}_{k}=\mathbf{K}_{k}^{0}$. The full algorithm is summarized in Algorithm 1. Note that we exactly recover the IEKF of [12] if we perform one iteration only.
@@ A-B Special case: noise-free measurements
In the limit case of noise-free measurements, i.e., $\mathbf{N}_{k}=\mathbf{0}$, the innovation covariance becomes $\mathbf{S}_{k}=\mathbf{H}_{k}\mathbf{P}_{k\mid k-1}\mathbf{H}_{k}^{T}$ and can cease to be invertible, in which case the Kalman gain $\mathbf{K}_{k}=\mathbf{P}_{k\mid k-1}\mathbf{H}_{k}^{T}\mathbf{S}_{k}^{-1}$ is undefined. This occurs when the perfectly observed directions overlap between subsequent updates for example. This issue is solved using the “noise-free limit gain” developed in our preliminary conference paper [34]:
@@ Appendix B Adaptation and extension to the LG-IterEKF
We stress that this result–a consequence of the Lie subalgebra structure related to the observation (7)–does not, at any rate, extend to the general LG-IterEKF presented in [5].

### v5

@@ IV Application of engineering interest
We propose using the IterIEKF to estimate the position, velocity, and orientation (extended pose) of the hook of a crane transporting a load, equipped with an IMU. Mounting an IMU on a crane hook for real-time sensor data transmission is technically very feasible nowadays [rauscher2018motion], and may open the door to new automation capabilities since it allows for feedback. Leveraging the Lie group $SE_{2}(3)$ introduced by invariant filtering theory in [barrau2016invariant], we represent the state as
@@
where $\mathbf{R}_{k}\in SO(3)$ is the rotation matrix between the IMU and inertial frames, and $\mathbf{v}_{k},\,\mathbf{p}_{k}\in\mathbb{R}^{3}$ are the IMU velocity and position in the inertial frame. The inertial frame is fixed at the crane cable attachment point with the $z$-axis oriented upward, and remains stationary. The hook and IMU frames are assumed perfectly aligned. See our preliminary conference paper [goffin2023invariant] for more details on this application.
@@ V Conclusion
We applied the IterIEKF to the problem of estimating the extended pose of a crane hook equipped with an IMU, where mechanical information is assimilated in the form of a noise-free pseudo-measurement. The IterIEKF demonstrated the best performance, closely followed by the adapted LG-IterEKF [bourmaud2016intrinsic], a fact explained by our present theory. The superiority of both algorithms is consistent with the theoretical guarantees we have proved, and the IEKF is outperformed in this context.
However, we recommend using the IterIEKF over the LG-IterEKF, because the IEKF possesses convergence properties (notably log-linearity of the error [barrau2016invariant]) and consistency properties that have never been proved for the LG-EKF. In particular, the unobservability consistency properties of the IEKF [barrau2018invariant, wu2017invariant, heo2018consistent] may not be inherited by the LG-IterEKF, although proving this fact is left for future research.
@@ A-A Standard case: noisy measurements
Akin to the conventional iterated EKF [bell1993iterated], we intend to solve (6) iteratively using the Gauss-Newton method.
Lemma 2 (from [bell1993iterated]).
@@
with $\mathbf{H}_{k}$ as in (10), and where the right Lie Jacobian [chirikjian2009stochastic] is defined by $\exp_{G}(\boldsymbol{\xi}+\boldsymbol{\delta})\approx\exp_{G}(\boldsymbol{\xi})\exp_{G}(\mathbf{\mathcal{J}}_{r}(\boldsymbol{\xi})\boldsymbol{\delta})$ neglecting terms of order $\|\boldsymbol{\delta}\|^{2}$, see (30) for an expression.
@@
with $\mathbf{H}_{k}=\mathbf{H}_{k}^{0}$ and $\mathbf{K}_{k}=\mathbf{K}_{k}^{0}$. The full algorithm is summarized in Algorithm 1. Note that we exactly recover the IEKF of [barrau2016invariant] if we perform one iteration only.
@@ A-B Special case: noise-free measurements
In the limit case of noise-free measurements, i.e., $\mathbf{N}_{k}=\mathbf{0}$, the innovation covariance becomes $\mathbf{S}_{k}=\mathbf{H}_{k}\mathbf{P}_{k\mid k-1}\mathbf{H}_{k}^{T}$ and can cease to be invertible, in which case the Kalman gain $\mathbf{K}_{k}=\mathbf{P}_{k\mid k-1}\mathbf{H}_{k}^{T}\mathbf{S}_{k}^{-1}$ is undefined. This occurs when the perfectly observed directions overlap between subsequent updates for example. This issue is solved using the “noise-free limit gain” developed in our preliminary conference paper [goffin2023invariant]:
@@ Appendix B Adaptation and extension to the LG-IterEKF
We stress that this result–a consequence of the Lie subalgebra structure related to the observation (7)–does not, at any rate, extend to the general LG-IterEKF presented in [bourmaud2016intrinsic].